# MIBoost: A Gradient Boosting Algorithm for Variable Selection After Multiple Imputation

### v3

@@ 1 Introduction
Multiple imputation is widely considered the gold standard for handling missing data. When the main interest lies in causal inference, coefficients and standard errors obtained from each imputed dataset can be easily aggregated using Rubin’s pooling rules [1]. However, performing variable selection for prediction purposes with multiple imputed datasets remains methodologically challenging, as the selected models can differ substantially across datasets. Many previously proposed methods for obtaining a single unified model are complex and insufficiently implemented in software packages. Simply selecting all variables chosen in any of the datasets often leads to overly complex models [2]; hence, many users resort to ad hoc solutions, such as including only variables selected in a minimum proportion of (e.g., 50%) of imputed datasets. While this heuristic may yield satisfactory results in some cases, it suffers from a strong dependence on an arbitrary threshold and can be unstable in small samples or in the presence of correlated predictors [3]. To address this problem, [4] proposed a framework that leverages the popular variable selection methods LASSO and elastic nets [5]. They propose a framework consisting of two different approaches, both of which minimize a single loss function, ultimately yielding a unified model. This framework is implemented in the R package miselect [6]. Another statistical-learning method often considered an alternative to LASSO and elastic nets is component-wise gradient boosting [7], which implicitly results in coefficient shrinkage and variable selection, too, making it well-suited for high-dimensional data. Whereas LASSO and elastic nets are limited to generalized linear models, boosting accommodates diverse loss functions and nonlinear effects, making it highly versatile. However, no method to date ensures uniform variable selection when applying component-wise gradient boosting to multiply imputed data. This paper extends the idea of defining a single loss function across imputations and proposes a novel algorithm for component-wise gradient boosting that enables consistent variable selection across all imputed datasets. This approach is implemented in the R package booami to enable accessibility for methodologically oriented as well as application-oriented researchers.
@@
  • We conduct a comprehensive simulation study assessing the predictive accuracy, variable-selection performance, and computational efficiency of MIBoost compared to existing boosting-based and imputation-based approaches.
@@ 2.1 Statistical and Component-Wise Gradient Boosting
$\hat{R}(\bm{\eta})=\sum_{i=1}^{n}\rho\big(y_{i},\eta(\bm{X}_{i})\big),$
[…]
$\bm{u}^{[t]}=-\nabla_{\bm{\eta}}\hat{R}(\bm{\eta}),\qquad\bm{u}^{[t]}\in\mathbb{R}^{n}.$
@@
these pseudo-residuals coincide with the ordinary residuals.
$\bm{u}=\bm{y}-\hat{\bm{\eta}},\qquad\bm{u}\in\mathbb{R}^{n}.$
@@
The additive predictor $\hat{\bm{\eta}}$ is a function of the design matrix $\bm{X}$, which is constructed from a set of $c$ raw covariates (e.g., age, sex). From these covariates, we derive a collection of $p$ candidate covariates ($c\leq p$), which form the columns of the design matrix,
@@
$\hat{h}_{\beta_{r}}^{[t]}=\arg\min_{h\in\mathcal{H}_{r}}\sum_{i=1}^{n}\big(u_{i}^{[t]}-h(x_{i,r})\big)^{2},\qquad\hat{h}_{\beta_{r}}^{[t]}:\mathbb{R}\to\mathbb{R}.$
@@
$\hat{\bm{h}}_{\beta_{r}}^{[t]}=\hat{\beta}_{0,r}^{[t]}\bm{1}+\hat{\beta}_{r}^{[t]}\bm{X}_{r},\qquad\hat{\bm{h}}_{\beta_{r}}^{[t]}\in\mathbb{R}^{n},$
@@
$\hat{\bm{\eta}}^{[t]}=\hat{\bm{\eta}}^{[t-1]}+\nu\hat{\bm{h}}_{\beta_{r^{*}}}^{[t]},\qquad\hat{\bm{\eta}}^{[t]}\in\mathbb{R}^{n}.$
@@ 3.1 The Algorithm MIBoost
(5) Post-processing: The above procedure is repeated for $t_{\mathrm{stop}}$ iterations. After boosting has concluded, the fitted base-learner functions are averaged across imputations to produce a single coherent final additive predictor. Although it is also possible not to pool the models and instead retain $M$ different models, the primary interest usually lies in obtaining a single model. Accordingly, we consider only the pooled model. Algorithm 3.2 summarizes the complete procedure. MIBoost is implemented in the function impu_boost of the R package booami [8], which also provides implementations of estimate averaging and selection-frequency thresholding - as defined in Subsection 2.2 - for direct comparison.
@@ 4.2 Results
Prediction accuracy was comparable across all evaluated methods (Table 1), with the boosting algorithms performing marginally better overall. In terms of variable-selection performance, all algorithms correctly identified the informative variables but also tended to include uninformative covariates. The model complexities and true negative proportions of MIBoost were similar to those of SaLASSO and SaENET. In contrast, gradient boosting with estimate averaging consistently tended to produced considerably more complex models and included almost two-thirds of the uninformative covariates.
@@ 6 Conclusion
Finally, since we did not apply standard LASSO or elastic nets but but rather their adaptive versions, we needed to define adaptive weight vectors for each run of SaLASSO or SaENET. To do so, we first fitted a preliminary stacked elastic net with equal weights (referred to as SENET in Du et al. [4]). We then extracted the estimated coefficients of all candidate covariates and used them to compute the corresponding adaptive weight vectors.

### v4

@@ 1 Introduction
Multiple imputation is widely considered the gold standard for handling missing data. When the main interest lies in causal inference, coefficients and standard errors obtained from each imputed dataset can be easily aggregated using Rubin’s pooling rules [1]. However, performing variable selection for prediction purposes with multiple imputed datasets remains methodologically challenging, as the selected models can differ substantially across datasets. Many previously proposed methods for obtaining a single unified model are complex and insufficiently implemented in software packages. Simply selecting all variables chosen in any of the datasets often leads to overly complex models [2]; hence, many users resort to ad hoc solutions, such as including only variables selected in a minimum proportion (e.g., 50%) of imputed datasets. While this heuristic may yield satisfactory results in some cases, it suffers from a strong dependence on an arbitrary threshold and can be unstable in small samples or in the presence of correlated predictors [3]. To address this problem, [4] proposed a framework that leverages the popular variable selection methods LASSO and elastic nets [5]. They propose a framework consisting of two different approaches, both of which minimize a single loss function, ultimately yielding a unified model. This framework is implemented in the R package miselect [6]. Another statistical-learning method often considered an alternative to LASSO and elastic nets is component-wise gradient boosting [7], which implicitly results in coefficient shrinkage and variable selection, too, making it well-suited for high-dimensional data. Whereas LASSO and elastic nets are limited to generalized linear models, boosting accommodates diverse loss functions and nonlinear effects, making it highly versatile. However, no method to date ensures uniform variable selection when applying component-wise gradient boosting to multiply imputed data. This paper extends the idea of defining a single loss function across imputations and proposes a novel algorithm for component-wise gradient boosting that enables consistent variable selection across all imputed datasets. This approach is implemented in the R package booami to enable accessibility for methodologically oriented as well as application-oriented researchers.
@@
  • We conduct a comprehensive simulation study assessing the predictive accuracy and variable-selection performance of MIBoost compared to existing boosting-based and imputation-based approaches.
@@ 2.1 Statistical and Component-Wise Gradient Boosting
$\hat{R}(\hat{\bm{\eta}}(\bm{X}))=\sum_{i=1}^{n}\rho\big(y_{i},\hat{\eta}(\bm{x}_{i})\big),$
which quantifies the discrepancy between the additive predictor $\hat{\eta}(\bm{X})$ and the observed responses $\bm{y}$. When evaluated at the fitted predictor $\hat{\eta}(\bm{X})$, this yields the empirical loss $\hat{R}(\hat{\eta}(\bm{X}))$. Boosting achieves minimization of this loss by iteratively descending $\hat{R}(\hat{\eta}(\bm{X}))$ along the direction of its negative gradient. Specifically, at each iteration, we compute
$\bm{u}^{[t]}=-\nabla_{\hat{\bm{\eta}}(\bm{X})}\hat{R}(\hat{\bm{\eta}}(\bm{X})),\qquad\bm{u}^{[t]}\in\mathbb{R}^{n},$
@@
these pseudo-residuals coincide with the ordinary residuals,
$\bm{u}=\bm{y}-\hat{\bm{\eta}}(\bm{X}),\qquad\bm{u}\in\mathbb{R}^{n}.$
@@
The additive predictor $\hat{\eta}(\bm{X})$ is a function of the design matrix $\bm{X}$, which is constructed from a set of $c$ raw covariates (e.g., age, sex). From these covariates, we derive a collection of $p$ candidate covariates ($c\leq p$), which form the columns of the design matrix,
@@
$\hat{h}_{\beta_{r}}^{[t]}=\arg\min_{h\in\mathcal{H}_{r}}\sum_{i=1}^{n}\big(u_{i}^{[t]}-h(x_{i,r})\big)^{2},\qquad\hat{h}_{\beta_{r}}^{[t]}:\mathbb{R}\to\mathbb{R},$
@@
$\hat{\bm{h}}_{\beta_{r}}^{[t]}(\bm{X}_{r})=\hat{\beta}_{0,r}^{[t]}\bm{1}+\hat{\beta}_{r}^{[t]}\bm{X}_{r},\qquad\hat{\bm{h}}_{\beta_{r}}^{[t]}(\bm{X}_{r})\in\mathbb{R}^{n},$
@@
$\hat{\bm{\eta}}^{[t]}(\bm{X})=\hat{\bm{\eta}}^{[t-1]}(\bm{X})+\nu\hat{\bm{h}}_{\beta_{r^{*}}}^{[t]}(\bm{X}_{r^{*}}),\qquad\hat{\bm{\eta}}^{[t]}(\bm{X})\in\mathbb{R}^{n}.$
@@ 3.1 The Algorithm MIBoost
(5) Post-processing: The above procedure is repeated for $t_{\mathrm{stop}}$ iterations. After boosting has concluded, the fitted base-learner functions are averaged across imputations to produce a single coherent final additive predictor. Although it is also possible not to pool the models and instead retain $M$ different models, the primary interest usually lies in obtaining a single model. Accordingly, we consider only the pooled model. Algorithm 3.2 summarizes the complete procedure. MIBoost is implemented in the function impu_boost of the R package booami [8], which also provides implementations of estimate averaging and selection-frequency thresholding—as defined in Subsection 2.2—for gradient boosting, enabling direct comparison.
@@ 4.2 Results
Prediction accuracy was comparable across all evaluated methods (Table 1), with the boosting algorithms performing marginally better overall. In terms of variable-selection performance, all algorithms correctly identified the informative variables but also tended to include uninformative covariates. The model complexities and true negative proportions of MIBoost were similar to those of SaLASSO and SaENET. In contrast, gradient boosting with estimate averaging consistently tended to produce considerably more complex models and included almost two-thirds of the uninformative covariates.
@@ 6 Conclusion
Finally, since we did not apply standard LASSO or elastic nets but rather their adaptive versions, we needed to define adaptive weight vectors for each run of SaLASSO or SaENET. To do so, we first fitted a preliminary stacked elastic net with equal weights (referred to as SENET in Du et al. [4]). We then extracted the estimated coefficients of all candidate covariates and used them to compute the corresponding adaptive weight vectors.